OpenStudy (peachy1):

math 3questions check them! MEDAL FAN

OpenStudy (jeansweets):

?

OpenStudy (peachy1):

@jeansweets

OpenStudy (error1603):

1m = 100cm

OpenStudy (error1603):

1000m = 1km

OpenStudy (peachy1):

ok

OpenStudy (error1603):

\[1200cm=\frac{ 1 }{ 100 }\]

OpenStudy (peachy1):

so thee answer is c

OpenStudy (error1603):

B

OpenStudy (error1603):

1 g = 1000 mg

OpenStudy (peachy1):

oh i see what i did rong thare

OpenStudy (error1603):

\[\frac{ 1g }{ ?1000mg }=\frac{ 1000mg }{1g}=1\]

OpenStudy (error1603):

Since 450,000 mg has mg in the numerator, and we want to cancel it, we multiply by the conversion factor that has mg in the denominator. \[450,000mg \times \frac{ 1g }{ 1000mg }=450\]

OpenStudy (peachy1):

ok

OpenStudy (error1603):

So you got it?

OpenStudy (peachy1):

yes thankyou

OpenStudy (error1603):

No problem

OpenStudy (error1603):

Glad to help